40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 20 23:26:10 2016, gautam, rana, Update, LSC, PSL and AUX-X temperatures changed 2016_01_20_ALS_OutOfLoop_1.pdf
    Reply  Thu Jan 21 18:34:04 2016, rana, Update, LSC, PSL and AUX-X temperatures changed 
       Reply  Fri Jan 22 01:56:01 2016, ericq, Update, LSC, Audio ALSX ALSX.wav
Message ID: 11942     Entry time: Thu Jan 21 18:34:04 2016     In reply to: 11940     Reply to this: 11943
Author: rana 
Type: Update 
Category: LSC 
Subject: PSL and AUX-X temperatures changed 

Is the black ref spectrum from this year or from May of 2015 or ?

I wonder if the noise is a bunch of fast spikes or if its a true broadband rumble. Maybe we can tell by looking at the analog DFD or PLL outputs?

ELOG V3.1.3-